OpenStudy (anonymous):

6[1-(-6)]8 its not 15

OpenStudy (anonymous):

Divide then multiply then add/subtract

OpenStudy (anonymous):

I would be interested in how you got the answer 15, if you are willing to post your work.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6[1-(-6)]+8 6[1+6]+8 6[7]+8 15

OpenStudy (anonymous):

i get 21

OpenStudy (anonymous):

you added a plus sign that was not in your original question which is correct?

OpenStudy (anonymous):

can you show work i have like 20 similar problems it would be great if you did

OpenStudy (anonymous):

I will work through it with you but first you must clarify if the question is \[6 \times (1-(-6) \times 8\] or\[6 \times (1-(-6)) +8\]

OpenStudy (anonymous):

Im having a problem because those brackets are not parenthesis

OpenStudy (anonymous):

I use the acrimony (I think thats what it is called) BEDMAS to remember my order of operations Brackets Exponents Division Multiplication Addition Subtraction I live in Canada and trying to remember as much as I can from previous schooling there should be no difference between the square brackets and the round brackets (parenthesis).

OpenStudy (anonymous):

tell me what you think 6 times 7 is

OpenStudy (anonymous):

i get 50 thats the answer so the brackets also mean multiply

OpenStudy (anonymous):

Does brackets were driving me crazy

OpenStudy (anonymous):

yes you assume multiplication when a number is written outside of the brackets like that.

OpenStudy (anonymous):

thank you much